OpenStudy (anonymous):

Solve for the variable. a + 5.1 = 10.9 A. 5.8 B. -5.8 C. 16 D. -16

OpenStudy (bahrom7893):

a = 10.9-5.1 a = 5.8

OpenStudy (anonymous):

a=10.9-5.1 =5.8